# DeWave: Discrete EEG Waves Encoding for Brain Dynamics to Text Translation

### 23 Dec 2023 v3

@@ Abstract
Figure 1: Overall illustration of translating EEG waves into text  [^2] This visualization still keeps teacher-forcing evaluation for fair comparison with previous methods.through quantised encoding.
@@ 1 Introduction
Decoding brain states into comprehensible representations has long been a focal point of research [19, 36, 8, 52]. Electroencephalogram (EEG) signals are particularly favored by researchers due to their non-invasive nature and ease of recording [29, 44]. Traditional EEG decoding techniques largely focus on classifying brain states into restricted categories like Motor Imaginary (MI) [35, 7], Emotion [18, 45], Robotic Control [42], and Gaming [26, 22]. However, these labels, bound to specific tasks, are insufficient for broad-based brain-computer communication. Consequently, there has been a surge of interest in brain-to-text (speech) translation in recent years. As the current trend leans towards large language models (LLM) [4, 11, 28] showcasing increasingly generalized intelligence capabilities, it becomes crucial to delve into ways of bridging the gap between brain signals and natural language representation. However, this area remains under-explored.
@@
The early work in brain-to-text translation [13, 1, 24, 39] relied on external event markers like handwriting or eye-tracking fixations to segment whole brain signals into fragmentary features. This methodology treated the task as word-level classification on a small, closed vocabulary set, with each time step analyzed individually. Both invasive [13, 40] (ECoG) and non-invasive [31] (EEG) brain signals have been used in these approaches. Notably, utilizing handwriting as event markers on invasive signals has led researchers [47] to achieve state-of-the-art (SOTA) recognition accuracy on closed-set character-level recognition. Wang [46] expanded the vocabulary size substantially (50,2652) and demonstrated the feasibility of open-vocabulary brain-to-text translation by employing pre-trained language models with word-level EEG features. However, limitations persist. The order of event markers, particularly eye-tracking fixations used to segment EEG waves into word-level features, may not match the natural word order in language  [^3] For example, the ZuCo Dataset collects data by simultaneously recording eye-tracking fixation and brain waves during reading tasks. However, the order of eye-fixations and spoken words may not always coincide.. Moreover, current methods do not have the capacity for direct text translation.
@@
In this paper, we present Discrete EEG Waves Encoding for Brain Dynamics to Text Translation (DeWave), a pioneering framework depicted in Figure 2. DeWave uses a vector quantized variational encoder to transform EEG waves into a discrete codex, linking EEG waves to tokens based on their proximity to codex book entries. This method offers two key advantages: it addresses significant distribution variances in EEG waves across individuals [27, 38], and rectifies order mismatches between raw wave sequences and text without eye-tracking markers. Our raw waves encoder is guided by both self-reconstruction and a contrastive supervision alignment between text embeddings and vectorized raw waves. To navigate the challenges of training with limited parallel data, DeWave leverages large-scale pre-trained language models [6, 34, 3], specifically employing BART [20], which combines BERT’s bidirectional context with GPT’s left-to-right decoder. Notably, our discrete codex aligns more closely with actual language tokens than continuous EEG features, serving as an interpretable bridge between EEG input and the language model.
@@ 3.3 EEG Vectorization
To obtain a semantically coherent codex, we introduce a cross-modality contrastive learning approach distinct from prevalent methods. Unlike CLIP [33, 21] that contrasts CLS embeddings between sample pairs within a mini-batch, our approach operates within a single EEG-text pair. We contrast the EEG-tokenized codex embeddings sequence $z_{q}$ with the text embeddings sequence $z_{t}$. Assuming the raw wave feature extractors can produce a token sequence in an organized chronological order, we treat the diagonal EEG codex and text word2vec encoding pairs as positive pairs within the sequences. All other pairs are considered negative. The model is trained to minimize the distance between embeddings of positive pairs and maximize that of the negative pairs. For a given EEG-text pair $(i,j)$, the loss is defined in Equation 4
@@ 5 Limitations
Despite DeWave’s enhancements in EEG-to-Text translation using a discrete codex and raw wave encoding, its accuracy remains far from real-life scenarios compared to traditional language-to-language translations. Also, to keep a fair comparison with EEG-to-Text, this paper uses a teacher-forcing setting when evaluating. This setting eliminates accumulation error and turns the sequence decoding task into a word-level classification task given the ground truth token from the previous step. This setting is much easier yet we think it is still valuable as it could suggest feature extraction quality while keeping a fair comparison.
Additionally, the experiments in this paper are restricted to public neural reading data, not fully align with the "silent speech" concept of direct thought translation from human brains. Instead, the current ZuCo dataset is collected by giving people reading stimuli. In that case, this is still “passively" trying to reconstruct the stimuli signal. This paper focuses on introducing Wav2Vec formation of feature extraction on raw EEG waves and introduces discrete codex as learnable representations for the EEG-to-Text translation domain. The scientific problem of which is a better way of doing the “silent speech" task remains unexplored for future works.
@@ 6 Conclusion
This paper presents DeWave, a framework for the recently proposed open-vocabulary EEG-to-Text translation task [46], introducing the concept of discrete codex encoding. This approach brings enhancement in corpus text relevancy metrics, such as BLEU and ROUGE. DeWave also expands the task setting to decode raw EEG waves without the assistance of eye fixation markers, offering a viable solution for the first time to enable direct translation. Despite these advancements, the quality of brain decoding results remains substantially low and still remains teacher forcing setting for fair comparison. The translation of thoughts directly from the brain is a valuable yet challenging endeavor that warrants significant continued efforts. Given the rapid advancement of Large Language Models, similar encoding methods that bridge brain activity with natural language deserve increased attention. In our future work, we will explore more reasonable settings that remove teacher forcing for both training and testing. Also include neural-feedback into this domain to increase the scientific nature of this domain.
@@ Acknowledgement
This work was supported in part by the Australian Research Council (ARC) under discovery grant DP210101093 and DP220100803, and the GrapheneX-UTS Human-Centric AI Centre sponsored by GrapheneX (2023-2031).

### 03 Jan 2024 v4

@@ Abstract
Figure 1: Overall illustration of translating EEG waves into text  [^2] This visualization still keeps teacher-forcing evaluation for a fair comparison with previous methods.through quantised encoding.
@@ 1 Introduction
Decoding brain states into comprehensible representations has long been a focal point of research [20, 38, 8, 54]. Electroencephalogram (EEG) signals are particularly favored by researchers due to their non-invasive nature and ease of recording [31, 46]. Traditional EEG decoding techniques largely focus on classifying brain states into restricted categories like Motor Imaginary (MI) [37, 7], Emotion [19, 47], Robotic Control [44, 55], and Gaming [28, 24]. However, these labels, bound to specific tasks, are insufficient for broad-based brain-computer communication. Consequently, there has been a surge of interest in brain-to-text (speech) translation in recent years. As the current trend leans towards large language models (LLM) [4, 12, 30] showcasing increasingly generalized intelligence capabilities, it becomes crucial to delve into ways of bridging the gap between brain signals and natural language representation. However, this area remains under-explored.
@@
The early work in brain-to-text translation [14, 1, 26, 41] relied on external event markers like handwriting or eye-tracking fixations to segment whole brain signals into fragmentary features. This methodology treated the task as word-level classification on a small, closed vocabulary set, with each time step analyzed individually. Both invasive [14, 42] (ECoG) and non-invasive [33] (EEG) brain signals have been used in these approaches. Notably, utilizing handwriting as event markers on invasive signals has led researchers [49] to achieve state-of-the-art (SOTA) recognition accuracy on closed-set character-level recognition. Wang [48] expanded the vocabulary size substantially and demonstrated the feasibility of open-vocabulary brain-to-text translation by employing pre-trained language models with word-level EEG features. However, limitations persist. The order of event markers, particularly eye-tracking fixations used to segment EEG waves into word-level features, may not match the natural word order in language  [^3] For example, the ZuCo Dataset collects data by simultaneously recording eye-tracking fixation and brain waves during reading tasks. However, the order of eye-fixations and spoken words may not always coincide.. Moreover, current methods do not have the capacity for direct text translation.
@@
In this paper, we present Discrete EEG Waves Encoding for Brain Dynamics to Text Translation (DeWave), a pioneering framework depicted in Figure 2. DeWave uses a vector quantized variational encoder to transform EEG waves into a discrete codex, linking EEG waves to tokens based on their proximity to codex book entries. This method offers two key advantages: it addresses significant distribution variances in EEG waves across individuals [29, 40, 9], and rectifies order mismatches between raw wave sequences and text without eye-tracking markers. Our raw waves encoder is guided by both self-reconstruction and a contrastive supervision alignment between text embeddings and vectorized raw waves. To navigate the challenges of training with limited parallel data, DeWave leverages large-scale pre-trained language models [6, 36, 3], specifically employing BART [21], which combines BERT’s bidirectional context with GPT’s left-to-right decoder. Notably, our discrete codex aligns more closely with actual language tokens than continuous EEG features, serving as an interpretable bridge between EEG input and the language model.
@@ 3.3 EEG Vectorization
To obtain a semantically coherent codex, we introduce a cross-modality contrastive learning approach distinct from prevalent methods. Unlike CLIP [35, 23, 22] that contrasts CLS embeddings between sample pairs within a mini-batch, our approach operates within a single EEG-text pair. We contrast the EEG-tokenized codex embeddings sequence $z_{q}$ with the text embeddings sequence $z_{t}$. Assuming the raw wave feature extractors can produce a token sequence in an organized chronological order, we treat the diagonal EEG codex and text word2vec encoding pairs as positive pairs within the sequences. All other pairs are considered negative. The model is trained to minimize the distance between embeddings of positive pairs and maximize that of the negative pairs. For a given EEG-text pair $(i,j)$, the loss is defined in Equation 4
@@ 5 Limitations
Despite DeWave’s enhancements in EEG-to-Text translation using a discrete codex and raw wave encoding, its accuracy remains far from real-life scenarios compared to traditional language-to-language translations. Also, to keep a fair comparison with EEG-to-Text, this paper uses a teacher-forcing setting in evaluation. This setting eliminates accumulation error and turns the sequence decoding task into a word-level classification task given the ground truth token from the previous step. This setting is relatively easier yet we think it is still valuable as it could suggest feature extraction quality while keeping a fair comparison.
Additionally, the experiments in this paper are restricted to public neural reading data, not fully aligning with the "silent speech" concept of direct thought translation from human brains. Instead, the current ZuCo dataset is collected by giving people reading stimuli. This paper focuses on introducing Wav2Vec formation of feature extraction on raw EEG waves and introduces discrete codex as learnable representations for the EEG-to-Text translation domain. One scientific problem in this domain is a better way of doing the “silent speech" task, which we are exploring as on-going research.
@@ 6 Conclusion
This paper presents DeWave, a framework for the recently proposed open-vocabulary EEG-to-Text translation task [48], introducing the concept of discrete codex encoding. This approach brings enhancement in corpus text relevancy metrics, such as BLEU and ROUGE. DeWave also expands the task to decode raw EEG waves without the assistance of eye fixation markers. Despite these advancements, the quality of brain decoding results remains substantially low and remains teacher forcing setting for fair comparison. The translation of thoughts directly from the brain is a valuable yet challenging endeavor that warrants significant continued efforts. In our ongoing work, we are exploring more reasonable settings that remove teacher forcing for both training and testing. We will also include the “neural-feedback" mechanism in this EEG-to-Text research to enhance the scientific value of this domain.
@@ Acknowledgement
This work was supported in part by the Australian Research Council (ARC) under discovery grants DP210101093 and DP220100803, and the GrapheneX-UTS Human-Centric AI Centre sponsored by GrapheneX (2023-2031).